# Research Paper

### v1

@@ Abstract
Over the last decade, the Australian Dark Energy (OzDES) collaboration has used Reverberation Mapping to measure the masses of high redshift supermassive black holes. Here we present the final review and analysis of this OzDES reverberation mapping campaign. These observations use years of photometric and spectroscopic observations of Active Galactic Nuclei (AGN) in the redshift range and bolometric luminosity range . Both photometry and spectra are observed in visible wavelengths, allowing for the physical scale of the AGN broad line region to be estimated from reverberations of the H, MgII and CIV emission lines. We successfully use reverberation mapping to constrain the masses of super-massive black holes, and combine with existing data to fit a power law to the lag-luminosity relation for the H and MgII lines with a scatter of dex, the tightest and most robust fit yet identified. We fit a similarly constrained relation for CIV, resolving a tension with the low luminosity literature AGN by accounting for selection effects. We also examine the impact of emission line width and luminosity (related to accretion rate) in reducing the scatter of these scaling relationships and find no significant improvement over the lag-only approach for any of the three lines. Using these relations, we further estimate the masses and accretion rates of AGN. We also use these relations to estimate the relative sizes of the H, MgII and CIV emitting regions, and find evidence that the MgII emission may occur further out than H. In short, we provide a comprehensive benchmark of high redshift AGN reverberation mapping at the close of this most recent generation of surveys, including light curves, time-delays, and the most reliable radius-luminosity relations to date.
@@
AGN in the local universe can have their masses constrained by the kinematics of their surrounding stars, either directly tracking orbits (e.g. Schodel_2002) or the bulk stellar velocity dispersion (i.e. the ‘ relation’ Ferrarese_2000; Gebhardt_2000), but such techniques become challenging at larger distances where entire galaxies appear as AGN-dominated point sources. The previous decade has also seen considerable progress in imaging two of the nearest SMBHs with interferometry at millimetre wavelengths (e.g. EHT_2019), and in using Very Large Telescope interferometry to potentially access finer angular resolutions (GravityPlus_2022). At higher redshifts, the limits on angular resolutions can also be surpassed with reverberation mapping (RM), a technique in which a long temporal baseline can substitute for a lack of spatial information (Blandford_McKee_1982; Peterson_1993).
Conceptually, reverberation mapping is straightforward: stochastic variations in the brightness of the AGN’s central engine drive variations in the brightness of gas in the surrounding broad-line region, but do so with a time delay associated with light travel time between the central engine and the BLR. Where we can distinguish signals from different regions, we can characterise the scale of this time delay and so infer the physical scale of the AGN (see Fig. 1 for a sketch of this process with simplified geometry). Knowledge of this geometric scale, coupled with a measure of the velocity and assumptions about kinematics, provides a means of constraining the mass of the SMBH. As such, reverberation mapping of AGN has become the primary means of constraining SMBH masses at any appreciable redshift, dominating over all other methods beyond (Cackett_2021).
For most of the history of reverberation mapping, data have been limited to studies of a hand-full of AGN at a time, predominantly at low redshift (e.g. Peterson_1999; CIV_Peterson_2005; Denney_2006), limited by the difficulty of continuous photometric and spectroscopic observations over the required time scales. This has changed with the advent of the first generation of ‘industrial scale surveys’, in which reverberation mapping is performed on hundreds of AGN at redshifts probing into the deep cosmological past. Over the last decade, surveys like the Australian Dark Energy Survey (OzDES; King_2015) and the Sloan Digital Sky Survey (SDSS; SDSS-Shen_2015) have made regular photometric and spectroscopic measurements of well over AGN in the redshift range , dramatically increasing the number of distant AGN observations with RM data to a statistically significant sample, and pushing our constraints on SMBH masses well into the high redshift domain.
@@
The paper is organised as follows. We discuss the physical principles of reverberation mapping in Section 2, including methods and limitations. In Section 3 we introduce the OzDES and external data sets used in this paper, and then in Section 4 we introduce the methodology used by OzDES, including lag measurement procedures and quality cuts, used to define this data set. We then present the full suite of recovered lags, SMBH masses and other results from the OzDES RM program in Section 5. In Section 6.1 we compare the relationship constraints of OzDES against those of SDSS and a wide range of other surveys for the H, MgII and CIV lines, identifying where they agree or are in tension, and use a combined data set to tightly constrain the slope, offset and scatter for all three lines. In Section 6.2 we confirm an absence of systematic biases from emission-line width or accretion rate. In Section 7 we derive a bolometric relation and use the different lag predictions for H, MgII and CIV as a probe of the relative scales of their emission regions. We then apply the updated relationships in Section 8 to derive single-epoch SMBH masses for AGN, and finally discuss the results of this paper, examining possible implications about the evolution of SMBH populations via our single epoch mass estimates, in Section 9.
@@ 2.1 Reverberation Mapping
The core of AGN reverberation mapping is based on arguments of the virialized orbit of the broad line region about the black hole, with the SMBH mass being correlated with the estimated radius of the BLR, , and its velocity dispersion as in equation 1. The complexities of the BLR’s unresolved geometry are captured in the broad uncertainties of the dimensionless ‘virial factor’, :111Mejia_Restrepo_2018 finds that virial factor may be anti-correlated with line width. In this work, we treat as being independent.
Though conceptually simple, each term in the numerator of equation 1 is non-trivial to measure for any given source. One hurdle is that the kinematics at play are dependent on the shape and dynamics of the BLR, the nature of which is as yet unresolved (Pancoast_2014a; Pancoast_2014b). This uncertainty is quantified by calibrating as a loosely constrained value from population analysis of nearby AGN, the masses of which are estimated through use of the relation (Woo_2015; Grier_2013b). Measurements of are made from the spectroscopic broadening of the reverberating emission lines, either by their full-width half-maximum (FWHM), or their dispersion (i.e. the second moment of the line profile about its peak), with different virial factors being tuned for each (see Section 3.3 for examples of RM papers using either technique).
High redshift AGN are beyond the distances at which we have been able to angularly resolve , and so RM instead uses the timescale at which light propagates through the AGN geometry to measure this size by proxy (see Figure 1). Variations in broad band photometric luminosity are dominated by the light emanating directly from the accretion disk ‘engine’, with these variations being echoed in the driven response in the emission lines of the BLR after some delay, , typically of order days to months. The BLR spectrum exhibits atomic recombination lines that can be easily distinguished from the broad AGN continuum emission, allowing us to identify this echo through spectroscopic observations even where our photometric observations are dominated by light from the accretion disk. With sufficiently tight cadence of photometric and spectroscopic observations, the delay (or lag) between driving continuum and driven emission line response, can be observed, allowing temporal resolution of observations to substitute for the lack of spatial resolution:
@@
For the nearest AGN (), RM is performed with H at and H at (e.g. HBETA_Bentz_2014). Beyond measuring the mean lag, some studies have even gone as far as using high signal to noise (SNR) measurements of hydrogen recombination lines to constrain the BLR structure and kinematics in further detail. In this “velocity resolved reverberation mapping” (e.g. Denney_2009; Grier_2013a), the reverberating line is treated as being affected by varying degrees of Doppler broadening from different velocities at different radii within the BLR, resulting in different parts of the line profile reverberating at different times.
At higher redshifts (), RM is performed with the MgII line at (e.g. MgII_Metzroth_2006). RM masses derived from MgII are complicated by iron contamination from the adjacent FeII emission lines, which can distort estimates of the line width unless accounted for (discussed in detail in Section 4).
At high redshift (), RM is performed with the CIV line at . CIV lag measurements have also been made by observing this line in the observer-frame UV range for low redshift and low luminosity sources (e.g. CIV_Rosa_2015; CIV_Peterson_2005; MgII_Metzroth_2006).
@@
The strong dependence of AGN lags on luminosity has given rise to the technique of stacking (Fine_2012_stacking; Fine_2013_stacking), in which sources of poor signal to noise but similar luminosity have their lag measurements combined. Through stacking, lag measurements can be made up in aggregate over many sources that would otherwise give poorly constrained results, such as done for the OzDES data in OzDES-Malik_2024.
@@
Here, is the cosine of the inclination of the AGN, taken to be as an average value for the quasar inclinations in which the BLR and accretion disk are both visible (Seyfert - 1 type AGN), while is the bolometric luminosity at in units of and is the RM-derived SMBH mass in units of solar masses. The Super-Eddington Accreting Mass Black Hole (SEAMBH) collaboration use as a benchmark for the separation between low accretion ”sub-Eddington” sources and the highly accreting ”super-Eddington” sources (Du_2015; HBETA_Du_2016).
@@ 2.2 The Aliasing Problem
Though the uncertainties in within a single source are dominated by the population variability in , lag-recovery presents significant issues with bias and contamination of mass estimates. Cosmological time-dilation means that distant AGN have longer timescales of variability (Lewis_2023), meaning that RM requires such sources to be observed over a baseline of multiple years to adequately capture variations in their light curves. For the ground-based observations of most distant RM-campaigns, such multi-year observations are necessarily impacted by a seasonal windowing function that imposes 6-month gaps in observations. These seasonal gaps give rise to the problem of ‘aliasing’ (see Fig. 2), which can yield spurious lag recoveries at yearly gaps (180 days, 420 days, etc.; OzDES-Penton_2021; OzDES-Malik_2022).
The dangers posed by aliasing are twofold; firstly, it degrades our ability to detect true lags that fall within seasonal gaps, and more dangerously it can lead to false positives arising within these same gaps (OzDES-Malik_2022). A common approach is to adopt weighting methods in combination with stringent selection criteria to remove sources with spurious lags. Such approaches reduce contamination from false positives, but at the cost of a drastically reduced sample, with typical acceptance ratios being of order in the OzDES sample. OzDES selections are based on agreement between competing methods (e.g. JAVELIN and ICCF; discussed in Section 4), the emergence of a single well constrained lag, and arguments about the physical reasonability of the overall lag posterior distribution (e.g. whether negative lags are properly excluded). More involved approaches rely on characterising the false positive rate (FPR) through simulation of sources (e.g. OzDES-Yu_2023; OzDES-Penton_2025). 222Recently, a new lag recovery code has been presented in the form of LITMUS,(McDougall_2025_litmus) which identifies false positives in a fully Bayesian way, but this method has not been used for any lag recoveries used in this paper.
@@ 3.1 OzDES
To complement the DES photometry, OzDES obtained monthly spectroscopy using the AAOmega spectrograph fed by the 2dF fibre positioner (Lewis_2002_2df) on the Anglo-Australian Telescope (AAT; for details see OZDES-DR2-Lidman_2020). The 2dF instrument has 400 optical fibres that can be positioned on targets within a 2-degree diameter field of view, and about a quarter of those were placed on AGN in each exposure. Using a spectral resolution of to and a wavelength range of to OzDES monitored a total of AGN over six years, gathering between and epochs on each AGN. OzDES typically observed for five months between August and January, perforce leaving large seasonal gaps in the time series data while the fields passed near to the sun. The spectra are publicly available from https://docs.datacentral.org.au/ozdes/overview/dr2/.
@@ 3.2 Light Curve & Spectral Calibration
Each spectral epoch is calibrated, both converting from photon counts to flux values and correcting for systematic distortions of the spectrum, by comparing the spectra to the , and photometric measurements. Synthetic photometric measurements are made by integrating over the product of the spectrum and band-pass, and a ratio taken between synthetic and measured brightness. Describing this factor as a function of wavelength, a quadratic is fit across the three measured ratios, acting as a measure of the relative throughput. The entire spectrum was warped to remove this bias (e.g. Fig. 3). After warping, spectra are co-added if they share the same night.
Once the spectrum is calibrated, line fluxes and widths are calculated by isolating a window of rest-frame wavelengths in which the reverberating line occurs. By linearly interpolating between the flux at the window boundaries, the local continuum is subtracted and the total flux, line dispersion and line width estimated. Uncertainties in this value are found by Monte Carlo variation of the window boundaries and the flux measurements within bounds of uncertainty.
For the MgII sources we add an additional step to account for the iron emission lines that flank and coincide with the MgII emission line, which, if not accounted for, can contaminate measurements of line width. In their analysis of the OzDES MgII lags, OzDES-Yu_2023 account for this by fitting iron spectral templates (Template_Tzuzuki_2006; Template_Salviander_2007; Template_Vestergaard_2001) in a Bayesian fashion using the python MCMC package emcee. Spectral fits for the emission lines were visually inspected and poorly fitted epochs were removed before lag-fitting with JAVELIN.

### v2

@@ Abstract
Over the last decade, the Australian Dark Energy (OzDES) collaboration has used Reverberation Mapping to measure the masses of high redshift supermassive black holes. Here we present the final review and analysis of this OzDES reverberation mapping campaign. These observations use years of photometric and spectroscopic observations of Active Galactic Nuclei (AGN) in the redshift range and bolometric luminosity range . Both photometry and spectra are observed in visible wavelengths, allowing for the physical scale of the AGN broad line region to be estimated from reverberations of the H, MgII and CIV emission lines. We successfully use reverberation mapping to constrain the masses of super-massive black holes, and combine with existing data to fit a power law to the lag-luminosity relation for the H and MgII lines with a scatter of dex, the tightest yet identified, fit specifically for consistency with high redshift AGN. We fit a similarly constrained relation for CIV, resolving a tension with the low luminosity literature AGN by accounting for selection effects arising from finite survey length. We also examine the impact of emission line width and luminosity (related to accretion rate) in reducing the scatter of these scaling relationships and find no significant improvement over the lag-only approach for any of the three lines. Using these relations, we further estimate the masses and accretion rates of AGN with single epoch methods. We also use these relations to estimate the relative sizes of the H, MgII and CIV emitting regions, and find evidence that the MgII emission may occur further out than H. In short, we provide a comprehensive benchmark of high redshift AGN reverberation mapping at the close of this most recent generation of surveys, including light curves, time-delays, and a set of significantly improved radius-luminosity relations for use with high-redshift populations.
@@
AGN in the local universe can have their masses constrained by the kinematics of their surrounding stars, either directly tracking orbits (e.g. Schodel_2002) or the bulk stellar velocity dispersion (i.e. the ‘ relation’ Ferrarese_2000; Gebhardt_2000), but such techniques become challenging at larger distances where entire galaxies appear as AGN-dominated point sources. The previous decade has also seen considerable progress in imaging two of the nearest SMBHs with interferometry at millimetre wavelengths (e.g. EHT_2019), and in using Very Large Telescope interferometry to potentially access finer angular resolutions (GravityPlus_2022). At higher redshifts, the limits on angular resolutions can also be surpassed with reverberation mapping (RM), a technique in which temporal coverage of variability can substitute for a lack of spatial information (Blandford_McKee_1982; Peterson_1993).
Conceptually, reverberation mapping is straightforward: stochastic variations in the brightness of the AGN’s central engine drive variations in the brightness of gas in the surrounding broad-line region, but do so with a time delay associated with light travel time between the central engine and the BLR. Where we can distinguish signals from different regions, we can characterise the scale of this time delay and so infer the physical scale of the AGN (see Figure 1 for a sketch of this process with simplified geometry). Knowledge of this geometric scale, coupled with a measure of the velocity and assumptions about kinematics, provides a means of constraining the mass of the SMBH. As such, reverberation mapping of AGN has become the primary means of constraining SMBH masses at any appreciable redshift, dominating over all other methods beyond (Cackett_2021).
For most of the history of reverberation mapping, data have been limited to studies of a hand-full of AGN at a time, predominantly at low redshift (e.g. Peterson_1999; CIV_Peterson_2005; Denney_2006), limited by the difficulty of continuous photometric and spectroscopic observations over the required time scales. This has changed with the advent of the first generation of ‘industrial scale surveys’, in which reverberation mapping is performed on hundreds of AGN at redshifts probing into the deep cosmological past. Over the last decade, surveys like the Australian Dark Energy Survey (OzDES; King_2015) and the Sloan Digital Sky Survey (SDSS; SDSS-Shen_2015) have made regular photometric and spectroscopic measurements of well over AGN in the redshift range , dramatically increasing the number of distant AGN observations with optical RM data to a statistically significant sample, and pushing our constraints on SMBH masses well into the high redshift domain.
@@
The paper is organised as follows. We discuss the physical principles of reverberation mapping in Section 2, including methods and limitations. In Section 3 we introduce the OzDES and external data sets used in this paper, and then in Section 4 we introduce the methodology used by OzDES, including lag measurement procedures and quality cuts, used to define this data set. We then present the full suite of recovered lags, SMBH masses and other results from the OzDES RM program in Section 5. In Section 6.1 we compare the relationship constraints of OzDES against those of SDSS and a wide range of other surveys for the H, MgII and CIV lines, identifying where they agree or are in tension in the plane, and use a combined data set to tightly constrain the slope, offset and scatter for all three lines for samples similar to OzDES and SDSS. In Section 6.2 we confirm an absence of systematic biases from emission-line width or accretion rate for these fits. In Section 7 we derive a bolometric relation and use the different lag predictions for H, MgII and CIV as a probe of the relative scales of their emission regions. We then apply the updated relationships in Section 8 to derive single-epoch SMBH masses for AGN, and finally discuss the results of this paper, examining possible implications about the evolution of SMBH populations via our single epoch mass estimates, in Section 9.
@@ 2.1 Reverberation Mapping
The core of AGN reverberation mapping is based on arguments of the virialised orbit of the broad line region about the black hole, with the SMBH mass being correlated with the estimated radius of the BLR, , and its velocity dispersion as in equation 1. The complexities of the BLR’s unresolved geometry are captured in the broad uncertainties of the dimensionless ‘virial factor’, :111Mejia_Restrepo_2018 finds that virial factor may be anti-correlated with line width. In this work, we treat as being independent.
Though conceptually simple, each term in the numerator of equation 1 is non-trivial to measure for any given source. One hurdle is that the kinematics at play are dependent on the shape and dynamics of the BLR, the nature of which is as yet unresolved (Pancoast_2014a; Pancoast_2014b). This uncertainty is quantified by calibrating the population average virial factor, from population analysis of nearby AGN, the masses of which are estimated through use of the relation (Woo_2015; Grier_2013b), or more recently through dynamical modelling via velocity resolved reverberation mapping (Villafana_2023_dynamical; Shen_2024_dynamical). The is only loosely constrained for any individual source, as there is an inherent scatter of dex about , and this dominates the uncertainty in BLR RM derived masses (Woo_2015; SDSS-Shen_2023).
Measurements of are made from the spectroscopic broadening of the reverberating emission lines, either by their full-width half-maximum (FWHM), or their dispersion (i.e. the second moment of the line profile about its peak), with different virial factors being tuned for each (see Section 3.3 for examples of RM papers using either technique).
High redshift AGN are beyond the distances at which we have been able to angularly resolve , and so BLR RM instead uses the timescale at which light propagates through the AGN geometry to measure this size by proxy (see Figure 1). Variations in broad band photometric luminosity are dominated by the light emanating directly from the accretion disk ‘engine’, with these variations being echoed in the driven response in the emission lines of the BLR after some delay, , typically of order days to months. The BLR spectrum exhibits atomic recombination lines that can be easily distinguished from the broad AGN continuum emission, allowing us to identify this echo through spectroscopic observations even where our photometric observations are dominated by light from the accretion disk. With sufficiently tight cadence of photometric and spectroscopic observations, the delay (or lag) between driving continuum and driven emission line response, can be observed, allowing temporal resolution of observations to substitute for the lack of spatial resolution:
@@
For the nearest AGN (), optical RM is performed with H at and H at (e.g. HBETA_Bentz_2014). Beyond measuring the mean lag, some studies have even gone as far as using high signal to noise (SNR) measurements of hydrogen recombination lines to constrain the BLR structure and kinematics in further detail. In this “velocity resolved reverberation mapping” (e.g. Denney_2009; Grier_2013a), the reverberating line is treated as being affected by varying degrees of Doppler broadening from different velocities at different radii within the BLR, resulting in different parts of the line profile reverberating at different times.
At higher redshifts (), optical RM is performed with the MgII line at (e.g. MgII_Metzroth_2006). optical RM masses derived from MgII are complicated by iron contamination from the adjacent FeII emission lines, which can distort estimates of the line width unless accounted for (discussed in detail in Section 4).
At high redshift (), optical RM is performed with the CIV line at . CIV lag measurements have also been made by observing this line in the observer-frame UV range for low redshift and low luminosity sources (e.g. CIV_Rosa_2015; CIV_Peterson_2005; MgII_Metzroth_2006).
@@
The strong dependence of AGN lags on luminosity has given rise to the technique of stacking (Fine_2012_stacking; Fine_2013_stacking; Li_2017_stacking), in which sources of poor signal to noise but similar luminosity have their lag measurements combined. Through stacking, lag measurements can be made up in aggregate over many sources that would otherwise give poorly constrained results, such as done for the OzDES data in OzDES-Malik_2024.
@@
Here, is the cosine of the inclination of the AGN, taken to be as an average value for the quasar inclinations in which the BLR and accretion disk are both visible (Seyfert - 1 type AGN), while is the bolometric luminosity at in units of and is the RM-derived SMBH mass in units of solar masses. The Super-Eddington Accreting Mass Black Hole (SEAMBH) collaboration use as a benchmark for the separation between low accretion “sub-Eddington” sources and the highly accreting “super-Eddington” sources (Du_2015; HBETA_Du_2016).
@@ 2.2 The Aliasing Problem
Though the uncertainties in within a single source are dominated by the dex population variability in , lag-recovery presents significant issues with bias and contamination of mass estimates. Cosmological time-dilation means that distant AGN have longer timescales of variability (Lewis_2023), meaning that RM requires such sources to be observed over a baseline of multiple years to adequately capture variations in their light curves. For the ground-based observations of most distant RM-campaigns, such multi-year observations are necessarily impacted by a seasonal windowing function that imposes 6-month gaps in observations. These seasonal gaps give rise to the problem of ‘aliasing’ (see Figure 2), which can yield spurious lag recoveries at yearly gaps (180 days, 420 days, etc.; OzDES-Penton_2021; OzDES-Malik_2022).
The dangers posed by aliasing are twofold; firstly, it degrades our ability to detect true lags that fall within seasonal gaps, and more dangerously it can lead to false positives arising within these same gaps (OzDES-Malik_2022). A common approach is to adopt weighting methods in combination with stringent selection criteria to remove sources with spurious lags. Such approaches reduce contamination from false positives, but at the cost of a drastically reduced sample, with typical acceptance ratios being of order in the OzDES sample. OzDES selections are based on agreement between competing methods (e.g. JAVELIN (JAVELIN-Zu_2010) and PyCCF (PyCCF-Mouyan_2018); discussed in detail in Section 4), the emergence of a single well constrained lag, and arguments about the physical reasonability of the overall lag posterior distribution (e.g. whether negative lags are properly excluded). More involved approaches rely on characterising the false positive rate (FPR) through simulation of sources (e.g. OzDES-Yu_2023; OzDES-Penton_2025). 222Recently, a new lag recovery code has been presented in the form of LITMUS,(McDougall_2025_LITMUS) which identifies false positives in a fully Bayesian way, but this method has not been used for any lag recoveries used in this paper.
@@ 3.1 OzDES
To complement the DES photometry, OzDES obtained monthly spectroscopy using the AAOmega spectrograph fed by the Two Degree Field (2dF) fibre positioner (Lewis_2002_2df) on the Anglo-Australian Telescope (AAT; for details see OZDES-DR2-Lidman_2020). The 2dF instrument has 400 optical fibres that can be positioned on targets within a 2-degree diameter field of view, and about a quarter of those were placed on AGN in each exposure. Using a spectral resolution of to and a wavelength range of to OzDES monitored a total of AGN over six years, gathering between and epochs on each AGN. OzDES typically observed for five months between August and January, perforce leaving large seasonal gaps in the time series data while the fields passed near to the sun. The spectra are publicly available from https://docs.datacentral.org.au/ozdes/overview/dr2/.
@@ 3.2 Light Curve & Spectral Calibration
Each spectral epoch is calibrated, both converting from photon counts to flux values and correcting for systematic distortions of the spectrum, by comparing the spectra to the , and photometric measurements. Synthetic photometric measurements are made by integrating over the product of the spectrum and band-pass, and a ratio taken between synthetic and measured brightness. Describing this factor as a function of wavelength, a quadratic is fit across the three measured ratios, acting as a measure of the relative throughput. The entire spectrum was warped to remove this bias (e.g. Figure 3). After warping, spectra are co-added if they share the same night.
Once the spectrum is calibrated, line fluxes and widths are calculated by isolating a window of rest-frame wavelengths in which the reverberating line occurs. By linearly interpolating between the flux at the window boundaries, the local continuum is subtracted and the total flux, line dispersion and line full width half maximum estimated. Uncertainties in this value are found by Monte Carlo variation of the window boundaries and the flux measurements within bounds of uncertainty.
Following the approach of HBETA_Du_2016, we apply a correction to the monochromatic luminosities of all H sources by way of the empirical scaling relationship provided by Equation 1 in shen_2011_stellarcontamination. A more thorough approach would be to decompose the spectra into host galaxy starlight and AGN activity, but is not crucial in this paper as the OzDES H sample’s small size means that it offers only weak constraining power to the parameters (see Section 6.1.1).
For the MgII sources we add an additional step to account for the iron emission lines that flank and coincide with the MgII emission line, which, if not accounted for, can contaminate measurements of line width. In their analysis of the OzDES MgII lags, OzDES-Yu_2023 account for this by fitting iron spectral templates (Template_Tzuzuki_2006; Template_Salviander_2007; Template_Vestergaard_2001) in a Bayesian fashion using the python Monte Carlo Markov Chain (MCMC) package emcee. Spectral fits for the emission lines were visually inspected and poorly fitted epochs were removed before lag-fitting with JAVELIN.